One week a business sold 40 scarves. White ones cost $4.95, and printed ones cost $7.95. In all, total sales of scarves were $282. How many of each kind were sold?
Answer by checkley77(12569)

P+W=40 OR P=40-W
4.95W+7.95(40-W)=282
4.95W+318-7.95W=282
-3W=282-318
-3W=-36
W=-36/3
W=12 WHITE SCARVES WERE SOLD.
40-12=28 PRINTED SCARVES WERE SOLD.
PROOF:
4.95*12+7.95*28=282
59.4+222.6=182
282=282